# Sparse Ellipsoidal Radial Basis Function Network for Point Cloud Surface Representation

### v2

@@ 4.6 Parameter Analysis
We analyze the key parameters used in SE-RBFNet, focusing on those that significantly affect model performance. The primary parameters include the threshold $\tau_{m}$ and the batch size $B_{s}$, both of which play a critical role in balancing speed, accuracy, and sparsity. The threshold $\tau_{m}$, introduced in Step 1 of Algorithm 2 and in Eq. 13, serves two purposes. First, during the basis function addition process, $\tau_{m}$ is used to assist in the identification of local extreme error points. Second, during optimization, once the max $\mathcal{L}_{2}$ error drops below $\tau_{m}$, an $\mathcal{L}_{1}$ regularization term is incorporated into the loss function to perform sparse optimization. We also conducted experiments with varying point sampling densities to assess reconstruction robustness under different sampling conditions. The other parameters are set as follows. To filter training points near the surface, we use two empirically set thresholds: $\tau_{1}=0.9$ and $\tau_{2}=1.1$. Adjusting these values controls the sampling region. Narrowing the band by setting $\tau_{1}\to 1^{-}$ and $\tau_{2}\to 1^{+}$ selects only points very close to the surface, which accelerates optimization but can reduce reconstruction accuracy. Conversely, widening the band increases computation but has little effect on accuracy. The deletion threshold for ERBFs is set to $\tau_{d}=0.01$, based on observations that most effective basis functions have weight magnitudes greater than 0.1. This setting helps eliminate negligible basis functions without harming performance. We cap the maximum number of iterations at 2000 due to our layered optimization strategy. Fewer iterations may lead to underfitting, while more iterations increase both runtime and the number of effective basis functions, as ERBFs are dynamically added or removed. Therefore, considering training efficiency, we limit the maximum number of iterations to 2000. The maximum octree depth is fixed at $l_{max}=10$, in line with baseline settings to ensure fair comparisons. Other parameters not explicitly mentioned follow the settings described in Section 4.1.
@@
(a) Armadillo Reconstruction Surface and Error Colormaps Under Different $\tau_{m}$
(b) Distance Error Distribution
(c) $\mathcal{L}_{2}$ Loss Change
(d) $\tau_{m}$ and ERBF Number
(e) Reconstruction Time
Figure 7: Effect of Threshold $\tau_{m}$ on Reconstruction Result, (a) Armadillo Reconstruction Surface and Error Colormaps Under Different $\tau_{m}$; (b) Distance Error Distribution between Reconstructed and Ground Truth Surface under Different $\tau_{m}$; (c) $\mathcal{L}_{2}$ Loss Change during Training under Different $\tau_{m}$; (d) Relationship between Final $\mathcal{L}_{2}$ Loss and the Number of Basis Functions under Different $\tau_{m}$; (e) Reconstruction Time Variation under Different $\tau_{m}$.
4.6.1 Impact of $\tau_{m}$ on Reconstruction Results
In Figure 7 we analyze the effect of different $\tau_{m}$ on reconstruction results. Figure 7a shows the reconstruction results and error colormaps under different $\tau_{m}$. It can be observed that as the threshold $\tau_{m}$ increases, surface details are gradually lost. The main reason for this is that $\tau_{m}$ acts as a switch that controls sparse optimization: once the max $\mathcal{L}_{2}$ error drops below $\tau_{m}$, sparse optimization is performed, leading to a reduction in the number of basis functions. Consequently, a larger $\tau_{m}$ triggers sparse optimization more frequently, which in turn leads to the model not being able to capture the details better. As shown in Figure 7a, some small bumps (protrusions) on the leg and some stripes on the abdomen. In Figure 7b, the histogram distribution describes the distance between the reconstructed point cloud and the ground truth surface for different $\tau_{m}$. It is clear that as the threshold value decreases, the histogram distribution becomes narrower and the peak higher. This indicates that the reconstruction error is concentrated in a smaller area, reflecting a higher accuracy. The error colormaps in Figure 7a also confirm this finding. Figure 7c shows the change in $\mathcal{L}_{2}$ loss for different $\tau_{m}$ during training. As $\tau_{m}$ decreases, the converged $\mathcal{L}_{2}$ loss during training becomes progressively smaller, indicating that a lower $\tau_{m}$ helps to fit the SDF more accurately to the actual surface. Figure 7d shows the relationship between the threshold $\tau_{m}$, the loss of $\mathcal{L}_{2}$ between the predicted SDF and the true SDF at the points of the octree grid, and the number of basis functions. The x-axis represents the final $\mathcal{L}_{2}$ loss corresponding to each threshold condition, while the y-axis represents the number of effective basis functions. It can be seen that as the threshold decreases, the final loss of $\mathcal{L}_{2}$ becomes smaller, the corresponding number of basis functions after optimization increases, and the required reconstruction time becomes longer (Figure 7e). Based on these observations, we conclude that setting the threshold $\tau_{m}$ to 0.0005 is the optimal balance. Compared to a smaller threshold of 0.00025, it achieves an almost equivalent surface reconstruction while reducing the number of basis functions by a third, thus significantly speeding up the reconstruction process. This trade-off ensures efficient reconstruction while maintaining a high level of accuracy.
(a) Reconstruction surface and basis function centers under Different $B_{s}$ (Utah_teapot_solid)
(b) Variation curves of CD, Time, $\mathcal{L}_{2}$ Loss and Effective Base Function Count under different $B_{s}$ (Utah_teapot_solid)
(c) Reconstruction surface and basis function centers under different $B_{s}$ (Horse)
@@ 4.6.1 Impact of $\tau_{m}$ on Reconstruction Results
(d) Variation curves of CD, Time, $\mathcal{L}_{2}$ Loss and Effective Base Function Count under different $B_{s}$ (Horse)
@@
Figure 8: Effect of batch size $B_{s}$ on Reconstruction Result. (a) and (c) show the reconstruction results and the distribution of the basis function centres for point clouds Utah_teapot_solid and Horse, respectively, under different batch sizes; (b) and (d) show the corresponding metric variation curves. From left to right, they show the variation of the chamfer distance under different batch sizes, the variation of the reconstruction time, the loss convergence during model training, and the change in the number of effective basis functions.
4.6.2 Impact of $B_{s}$ on reconstruction results
We observe that different batch sizes $B_{s}$ also significantly affect the reconstruction results. In Figs. 8, we analyze the effect of different batch sizes on the reconstruction performance. We selected two models from the famous dataset for testing, namely Utah_teapot_solid and Horse. Figs. 8a and 8c display the reconstruction results and center distributions of the basis function under different $B_{s}$. As $B_{s}$ increases, the number of effective basis functions gradually increases. This trend is confirmed in the fourth subfigure of both Figs. 8b and 8d, where the number of effective basis functions for the Utah_teapot_solid increases from 226 (at $B_{s}$ = 2000) to 1119 (at $B_{s}$ = 20,000), while for the Horse, it increases from 689 (with $B_{s}$ = 2000) to 1935 (with $B_{s}$ = 20,000). This phenomenon can be attributed to the increased instability of weight induced by the regularization of $\mathcal{L}_{1}$ used during training. Smaller batch sizes exacerbate weight fluctuations, causing certain basis functions to oscillate frequently during training and, ultimately, leading to pruning. In contrast, larger batch sizes result in more stable gradients and consistent weight updates, making the sparsification mechanism less effective at pruning a sufficient number of basis functions.
Figs. 8b and 8d further illustrate the corresponding metric variation curves. From left to right, the subfigures correspond to the variation of CD in different batch sizes, changes in reconstruction time, loss convergence during model training, and the number of effective basis functions. It is evident that as $B_{s}$ increases, CD decreases, primarily due to the increase in effective basis functions, which allows the model to capture more surface geometry details. The reconstruction time initially decreases rapidly and then increases gradually. The initial decrease is primarily due to the efficient parallel data processing enabled by larger batch sizes on the GPU. The subsequent increase is attributed to the increase in the number of basis functions, which increases the computational cost per epoch, thereby increasing the total training time. The third subfigure shows the loss convergence over the final 150 epochs. As the batch size increases, the model’s fitting ability improves, leading to a lower convergence loss. Based on these observations, we set $B_{s}=10000$ for the final experiments.
4.6.3 Impact of Point Cloud Sampling Density
(a) Reconstruction HD with Different Sampling Densities (10K, 20K, 30K, 40K Points)
(b) Reconstruction CD with Different Sampling Densities (10K, 20K, 30K, 40K Points)
(c) Reconstruction CS with Different Sampling Densities (10K, 20K, 30K, 40K Points)
(d) Reconstruction Time with Different Sampling Densities (10K, 20K, 30K, 40K Points)
(e) Number of Effective Basis Functions with Different Sampling Densities (10K, 20K, 30K, 40K Points)
@@ 4.6.3 Impact of Point Cloud Sampling Density
Figure 9: Effect of Sampling Densities on Reconstruction Result. The results for the ABC, Famous, MeshSeg and Thingi10K datasets are shown in the four columns from left to right.
The sampling density of the point cloud also influences the reconstruction result. A higher sampling density allows the point cloud to more comprehensively cover the target surface, thereby capturing finer geometric details. Conversely, sparse sampling increases the distance between adjacent points, which makes the reconstruction more susceptible to interpolation errors and detail loss. To systematically assess this effect, we conducted experiments with varying numbers of sampling points, and the corresponding results are presented in Figure 9, where the red line denotes the performance of our method. Figures 9a, 9b, and 9c present the evaluation results of reconstruction quality under different sampling densities. Overall, as the number of sampling points increases, our method exhibits a clear and consistent trend: HD and CD values decrease, while CS scores increase, indicating that the reconstruction quality improves with higher sampling density. This observation is consistent with the general trend discussed earlier. In contrast, SparseRBF fails to show a similarly stable trend under most conditions, suggesting its limited adaptability to varying sampling densities and lower robustness in maintaining reconstruction quality. Figure 9d illustrates the reconstruction time as a function of sampling density. While SparseRBF exhibits a significant increase in computational time as the number of points grows, our method maintains a much more stable trend, with only a moderate increase, further demonstrating its scalability and efficiency for large-scale point cloud reconstruction. Finally, Figure 9e shows the variation in the number of basis functions as the sampling density increases. Although the total number of basis functions increases relatively with increasing sampling density, our method consistently requires approximately 50% fewer basis functions than SparseRBF, highlighting its advantages in terms of storage efficiency and transmission for surface representations.
@@ 5 Conclusion and Future Work
This paper introduces SE-RBFNet, a sparse ellipsoidal radial basis function network designed for efficient and compact implicit surface representation. Specifically, the point cloud surface representation problem is reformulated as a sparse nonlinear optimization task, where the SDF for implicit surface representation of the target surface is approximated using a set of ERBFs. To address the trade-off between accuracy and model complexity, we propose a dynamic multi-objective optimization framework, combined with an octree-based coarse-to-fine hierarchical strategy and a nearest-neighbor filtering mechanism, to accelerate training and reduce redundancy. Furthermore, an adaptive basis function adjustment mechanism is introduced, which iteratively selects new centers from high-error regions and dynamically updates the ellipsoidal parameters, enabling the network to represent complex surfaces with significantly fewer basis functions. Extensive experiments demonstrate that SE-RBFNet achieves high-fidelity implicit surface representations with a less number of RBF bases, making it highly suitable for applications that balance efficiency with storage and transmission requirements. In future work, we will further explore adaptive strategies for basis function selection and parameter initialization, as well as investigate hybrid architectures that combine the advantages of sparse geometric modeling and neural implicit representations. Additionally, we aim to explore the application of sparse representations to other challenges, such as surface matching and other related problems, where the compact and efficient nature of sparse models can offer significant benefits.
@@ Appendix A Complete Gradient Analysis for SE-RBFNet
SE-RBFNet optimises surface representation problems using the following loss function:
@@
$\Psi(\bm{v}_{i},\bm{\widetilde{\Theta}})$ is the SE-RBFNet output for the input $\bm{v}_{i}$, $t_{i}$ is the corresponding ground-truth SDF value, and $\bm{\widetilde{\Theta}}$ denotes the set of all SE-RBFNet parameters. Only the points corresponding to the indices in $id_{all}$ are used for loss calculation.

### v4

@@ 4.5 Parameter Analysis
We analyze the key parameters used in SE-RBFNet, focusing on those that significantly affect model performance. The primary parameters include the threshold $\tau_{m}$ and the batch size $B_{s}$, both of which play a critical role in balancing speed, accuracy, and sparsity. The threshold $\tau_{m}$, introduced in Step 1 of Algorithm 2 and in Eq. 13, serves two purposes. First, during the basis function addition process, $\tau_{m}$ is used to assist in the identification of local extreme error points. Second, during optimization, once the max absolute error drops below $\tau_{m}$, the $\mathcal{L}_{1}$ regularization term is incorporated into the loss function to perform sparse optimization. We also conducted experiments using area-weighted uniform sampling on the mesh surface with varying numbers of sampled points to evaluate the robustness of the implicit surface approximation. Other parameters not explicitly mentioned follow the settings described in Section 4.1.
4.5.1 Impact of $\tau_{m}$ on Approximation Results
@@
(a) Armadillo Surface Approximation Results and Error Colormaps under Different $\tau_{m}$
@@
(b) Variation of CD, Time, $\mathcal{L}_{2}$ Loss and Basis Function Count under Different $\tau_{m}$
@@
Figure 6: Effect of $\tau_{m}$ on implicit surface approximation results, (a) Armadillo implicit surface approximation results and error colormaps under different $\tau_{m}$; (b) shows the corresponding metric variation curves. From left to right, these curves represent the variation of the chamfer distance with different $\tau_{m}$, computational times, loss convergence during training, and the number of effective basis functions.
In Figure 6, we analyze the effect of different $\tau_{m}$ on approximation results. Figure 6a shows the implicit surfaces extracted from the approximated SDF and corresponding error colormaps under various $\tau_{m}$. As $\tau_{m}$ increases, finer surface details tend to be lost. This occurs because $\tau_{m}$ controls the sparse optimization trigger: when the maximum absolute error falls below $\tau_{m}$ (see Eq. 13), sparse optimization prunes basis functions, reducing their number. A larger $\tau_{m}$ thus activates pruning earlier and more frequently, potentially causing insufficient detail capture. For example, small bumps on the leg and stripes on the abdomen become less pronounced at higher $\tau_{m}$. Figure 6b (leftmost) presents the CD between the implicit surfaces extracted from the approximated SDF and the ground-truth surface. Smaller values of $\tau_{m}$ lead to lower CD, indicating higher approximation accuracy. The error colormaps in Figure 6a also confirm this finding. Figure 6b (third from left) illustrates the $\mathcal{L}_{2}$ loss during training under different $\tau_{m}$ values, where smaller thresholds lead to lower converged losses, reflecting better SDF fitting. Figure 6b (rightmost) depicts the relationship between $\tau_{m}$, final $\mathcal{L}_{2}$ loss, and the number of basis functions after optimization. Decreasing $\tau_{m}$ results in lower loss but requires more basis functions and longer training time, as shown in Figure 6b (second from left). Based on these observations, $\tau_{m}=0.02$ strikes an optimal balance between approximation accuracy and efficiency: it achieves nearly the same accuracy as $\tau_{m}=0.015$ while using roughly half the number of basis functions and significantly reducing computation time.
4.5.2 Impact of $B_{s}$ on Approximation Results
(a) Implicit Surface Approximation Results and Distributions of Basis Function Centers under Different $B_{s}$ (Utah_teapot_solid)
(b) Variation of CD, Time, $\mathcal{L}_{2}$ Loss and Basis Function Count under Different $B_{s}$
Figure 7: Effect of batch size $B_{s}$ on implicit surface approximation results. (a) shows the implicit surface approximation results and the distribution of the basis function centers for the Utah_teapot_solid surface under different batch sizes; (b) shows the corresponding metric curves, from left to right: chamfer distance, training time, loss convergence, and the number of effective basis functions.
Batch size $B_{s}$ also influences the approximation quality. Figure 7 shows results on the Utah_teapot_solid model from the Famous dataset. Figure 7a displays the approximated implicit surfaces and basis function center distributions for varying $B_{s}$. Increasing $B_{s}$ leads to more effective basis functions: as shown in the rightmost plot of Figure 7b, the number of basis functions increases from 175 at $B_{s}=2000$ to 757 at $B_{s}=20000$. This is attributed to training stability: smaller batch sizes induce weight fluctuations under $\mathcal{L}_{1}$ regularization, causing basis functions to oscillate and be pruned, while larger batch sizes yield steadier updates and less pruning. The leftmost and third plots of Figure 7b indicate that the CD decreases and loss converges lower as $B_{s}$ increases, reflecting improved fitting accuracy due to more basis functions. Furthermore, the second plot of Figure 7b shows that training time first decreases with $B_{s}$ owing to efficient GPU parallelization, but then increases again as excessively large batch sizes induce a greater number of basis functions, leading to longer training times. Therefore, we select $B_{s}=10000$ as a compromise between approximation accuracy and computational cost for this test case.
4.5.3 Impact of the Number of Sampled Points
(a) CD with Different Numbers of Sampled Points (10K, 20K, 30K, 40K)
(b) CS with Different Numbers of Sampled Points (10K, 20K, 30K, 40K)
(c) Time with Different Numbers of Sampled Points (10K, 20K, 30K, 40K)
(d) Basis Functions with Different Numbers of Sampled Points (10K, 20K, 30K, 40K)
@@ 4.5.3 Impact of the Number of Sampled Points
Figure 8: Effect of the number of sampled points on implicit surface approximation results. Average results across the ABC, Famous, MeshSeg, and Thingi10K datasets for varying numbers of sampled points are shown in the four columns from left to right.
The number of sampled points directly affects the quality of SDF approximation, which in turn impacts the resulting implicit surface representation. More sampled points better cover the surface, capturing finer geometric details; sparser sampling increases interpolation difficulty and detail loss. Figure 8 summarizes the average performance across the ABC, Famous, MeshSeg, and Thingi10K datasets for different numbers of sampled points, with our method shown by the red curve. As illustrated in Figures 8a and 8b, both our method and SparseRBF show that increasing the number of sampled points leads to lower CD and higher CS, indicating improved implicit surface quality. Nevertheless, our method consistently achieves better metric results across different numbers of sampled points. Figure 8c illustrates that our method scales more gracefully with the number of sampled points, incurring only moderate increases in approximation time, while SparseRBF’s time grows sharply. Figure 8d indicates that the number of basis functions in our method increases slightly with the number of sampled points, in contrast to SparseRBF, where it rises as the number of sampled points grows. Moreover, our method consistently requires significantly fewer basis functions, highlighting efficiency advantages in the storage and transmission of implicit surface representations.
@@ 5 Conclusion and Future Work
This paper introduces SE-RBFNet, a sparse ellipsoidal radial basis function network designed for efficient and compact approximation of SDFs of implicit surfaces. Unlike methods that generate SDFs, our approach focuses on the sparse approximation of precomputed SDF samples, which can be derived from point clouds, triangular meshes, analytical SDFs, or neural implicit models. Specifically, the implicit surface representation problem is reformulated as a sparse nonlinear optimization task: given SDF values sampled on a spatial grid, SE-RBFNet approximates them using a small set of ellipsoidal RBFs, achieving a sparse representation that preserves the zero-level set of complex geometries while significantly reducing the number of parameters. To balance accuracy and model complexity, we propose a dynamic multi-objective optimization framework, combined with a coarse-to-fine hierarchical strategy and a nearest-neighbor filtering mechanism, to accelerate training and reduce redundancy. Furthermore, an adaptive basis function adjustment mechanism is introduced, which iteratively selects new centers from high-error regions and dynamically updates the ellipsoidal parameters, enabling the network to represent complex surfaces with significantly fewer basis functions. Extensive experiments demonstrate that SE-RBFNet can accurately approximate complex SDFs with substantially fewer parameters, achieving high-fidelity implicit surface representation while reducing storage and computational costs.
In future work, we will further explore adaptive strategies for basis function selection and parameter initialization, as well as investigate hybrid architectures that combine the advantages of sparse geometric modeling and neural implicit representations. Additionally, we aim to explore the application of sparse representations to other challenges, such as surface matching and other related problems, where the compact and efficient nature of sparse models can offer significant benefits.
@@ Appendix A Complete Gradient Analysis for SE-RBFNet
SE-RBFNet addresses the sparse approximation of SDFs for implicit surfaces by optimizing the following loss function:
@@
[…]